OpenStudy (anonymous):

Solve algebraically. Formula below:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin (4x) \over \sin(3x)\]

OpenStudy (anonymous):

you can use l'hopital's rule. it says when you have a case where you get 0/0 or infinity/infinity, then you can take the first derivative of the top function and divide it by the first derivative of the bottom function.

OpenStudy (anonymous):

Divide by the highest coff. which is x so the answer will be 4/3

OpenStudy (anonymous):

This was on the test yesterday. We don't know l'hospital's rule yet :( Previous examples showed manipulating the problem so that you can use lim x -> 0 \[\sin (x) \over x\] = 1

OpenStudy (anonymous):

@glorydotcom : Here is the steps: \[\huge \lim_{x \rightarrow 0} \frac{\sin4x}{\sin3x}\] Divide by the highest coff. of x which is x^1. \[\huge \lim_{x \rightarrow 0} \frac{\frac{(\sin4x)}{(x)}}{\frac{(\sin3x)}{(x)}}\] = \[\huge \frac{4}{3}\]

OpenStudy (anonymous):

l'hospital's rule's the easiest way to do it.. and the only way i can think up right now.. >< so the derivative of sin(4x) = 4cos(x) and similarly, sin(3x) = 3cos(x) when x-> 0, you get 4(1)/3(1) to get 4/3. oh. what @Eyad did above is almost correct. you can do that, but sin(3x)/x doesn't simplify to 3 like that. the answer is 3, but it's not because the x's have cancelled.

OpenStudy (anonymous):

@nuk64 :actually i didn't say anthing about canceling ,Did I ?

OpenStudy (anonymous):

so for sin(3x)/x, what you do is multiply the top and bottom by 3. so you do sin(3x)*3/x*3 then you get 3sin(3x)/3x which is the same property - sin(x)/x which goes to 1. so you have 3*1 which equals 3. :) same goes for the sin(4x). oh. true. sorry @Eyad ><

OpenStudy (anonymous):

@nuk64 :Nvm :) and btw good explanation :)

OpenStudy (anonymous):

Thanks guys!